Which equation has a slope
m = 1/2
OPTIONS

A) y = 2x + 1

B) y = -2x - 1/2

C) y = - 1/2 x + 1/2

D) y = 1/2 x + 3


Answer by jim_thompson5910(35256)

The answer is y = 1/2x + 3 because this equation is in the form y = mx+b

m is the slope.

In this case, the slope is m = 1/2